I would like assitance with these problems if available and would appreciate you.
Describe the transformations onthe following graph of f(x)= e^x. State the placement of the horizontal asymptote and y-intercept after the transformation.
a. g(x)=e^x -2
Description of transformation:
Horizontal asymptote:
y-intercept in (x,y) form
an the same b. h(x)=-e^x

3) Describe the tranformations onthe following graph of f(x)=log(x)
g(x)=log(x-3)
Description of transformation:
Vertical asymptote:
x-intercept in (x,y) form:
b) g(x)=log(-x)
Description of transformation:
Vertical asymptote:
x-intercept in (x.y) form
I just get confuse on the formulas

Answer by stanbon(75887)

Describe the transformations onthe following graph of f(x)= e^x. State the placement of the horizontal asymptote and y-intercept after the transformation.
a. g(x)=e^x -2
Description of transformation:
Horizontal asymptote: y = -2
--------
y-intercept in (x,y) form: Let x = 0 then y = 1-2=-1
---------------------------
and the same b. h(x)=-e^x
Horizontal asymptote: y = 0
y-intercept: Let x=0, y = -1
---------------------------
3) Describe the tranformations onthe following graph of f(x)=log(x)
g(x)=log(x-3)
Description of transformation:
Shift all points of f(x) three units to the right.
---------------
Vertical asymptote: x = 3
----------
x-intercept in (x,y) form:
Let y = 0 : log(x-3)=0 ; x-3 = 10^0 ; x-3 = 1; x = 4
----------------------
b) g(x)=log(-x)
Description of transformation:
Reflect f(x) in the y axis
-------------
Vertical asymptote:
x=0
----------
x-intercept in (x.y) form
Let y=0; 0 = log(-x) ; -x = 10^0; -x = 1 ; x= -1
----------------
===================
Cheers,
Stan H.